OpenStudy (anonymous):

A clock with a radius of 15 cm has an 11 minute hand on a 7cm hour hand. How far apart are the tips of the hands at 3:30. Hint: you use the sine or cosine law to solve this

OpenStudy (radar):

It would appear that the angle included between the hour and the minute hand at 3:30 would be 75 degrees 2.5/12 X (360) degrees = 75 degrees. You now have the included angle of 75 degrees, with one side being 11 the other being 7, you can find the third side which is the desired answer

OpenStudy (radar):

\[c ^{2}=7^{2}+11^{2}-2(7)(11)\cos 75^{0}\]

OpenStudy (radar):

\[c ^{2}=49+121-(2)(7)(11)(.2588)\] \[c ^{2}=170-39.8552\] \[c ^{2}=130.1448\] \[c=11.408 cm\]

OpenStudy (radar):

Check for errors lol